OpenStudy (chrisplusian):

Need help understanding why my answer is wrong on a direct comparison test.

OpenStudy (chrisplusian):

\[\sum_{0}^{\infty}\frac{ 1 }{ n! }\]

OpenStudy (anonymous):

I wonder what you could compare this to...

OpenStudy (chrisplusian):

I saw the n factorial terms as\[\frac{ 1 }{ n }+\frac{ 1 }{ n-1 }+\frac{ 1 }{ n-2 }+.....\frac{ 1 }{ 2 }+\frac{ 1 }{ 1 }\] then compared that to \[\frac{ 1 }{ n }\]which is a divergent p-series. Is it wrong to say that \[\frac{ 1 }{ n } \le \frac{ 1 }{ n! }\]????

OpenStudy (anonymous):

Whoa there, where did you get \[\frac{ 1 }{ n }+\frac{ 1 }{ n-1 }+\frac{ 1 }{ n-2 }+.....\frac{ 1 }{ 2 }+\frac{ 1 }{ 1 }\]?

OpenStudy (chrisplusian):

I goofed I am thinking multiplication

OpenStudy (anonymous):

So... you found your error? :D

OpenStudy (chrisplusian):

so maybe i Have the concept confused. When I am using the direct comparison test am I comparing the nth partial sums? Or am I comparing the sequence?

OpenStudy (anonymous):

The sequence :) Each individual term of the sequence, you compare

OpenStudy (chrisplusian):

so then the 1/n! would be\[\frac{ 1 }{ 0! }+\frac{ 1 }{ 1! }+\frac{ 1 }{ 2! }\] etc?

OpenStudy (anonymous):

Yes :D

OpenStudy (chrisplusian):

and I need to find a sequence that diverges that is less than this sequence or one that converges that is greater than this sequence?

OpenStudy (anonymous):

Yup.

OpenStudy (chrisplusian):

Yikes!!

OpenStudy (anonymous):

It doesn't have to be greater all the time. Just that at some point, it will always be greater/smaller

OpenStudy (chrisplusian):

right but I thought factorials grew faster than exponentials?

OpenStudy (anonymous):

In this case, we could use this (convergent) geometric series \[\large \frac1{2^n}\] Now, for n = 0,1,2, and 3 \[\Large \frac1{n!}>\frac1{2^n}\] But for n = 4 onwards \[\large \frac1{n!}<\frac1{2^n}\] So you can use the comparison test there ^.^

OpenStudy (chrisplusian):

I am a little confused, I thought somewhere along the line someone said that factorials grew faster than exponentials.

OpenStudy (anonymous):

of course, that's why when factorials are in the DENOMINATOR, they become smaller than when you have EXPONENTIALS at your denominator :)

OpenStudy (chrisplusian):

Duh LOL I was just testing you see if you remembered lol

OpenStudy (anonymous):

Is that so? -.-

OpenStudy (chrisplusian):

Thanks that helps. I was just having a slow moment